OpenStudy (anonymous):

rewrite the expression as an algebraic expression in x sin(2sin^-1(2x))

OpenStudy (beccaboo333):

Welcome to OpenStudy @secnorc what exactly are you having trouble on? Do you understand how to do this or not at all/

OpenStudy (anonymous):

I've solved some like cos(tan^-1(3x)), but this one I am not getting

OpenStudy (anonymous):

Are you able to walk me throught it?

OpenStudy (raden):

use these identitties : sin(2x) = 2sin(x)cos(x) and sin^2 (x) + cos^2 (x) = 1

OpenStudy (anonymous):

I'm still not understanding where to go

OpenStudy (raden):

sin(2sin^-1(2x)), in parantheses is double angles right ? so, it can rewitten as : sin(2sin^-1(2x)) = 2 sin(sin^-1(2x)) cos sin(sin^-1(2x)), agree ?

OpenStudy (anonymous):

okay, yes

OpenStudy (raden):

2 sin(sin^-1(2x)) cos sin(sin^-1(2x)) lets do slowly ;) obvious, for sin(sin^-1(2x)) = 2x, right ?

OpenStudy (anonymous):

Right

OpenStudy (raden):

ok, we get 2 sin(sin^-1(2x)) cos sin(sin^-1(2x)) = 2 (2x) cos sin(sin^-1(2x)) = (4x) cos sin(sin^-1(2x)) now, still be a problem for cos sin(sin^-1(2x)), right ?

OpenStudy (anonymous):

yeah, would it come out being (4x)cos(2x) next?

OpenStudy (raden):

oppps,, typo there that should like this : 2 sin(sin^-1(2x)) cos(sin^-1(2x)) = 2 (2x) cos(sin^-1(2x)) = 4x cos(sin^-1(2x)) now the problem is for evaluate the cos(sin^-1(2x))

OpenStudy (anonymous):

is it sqrt(4x^2+1)/1?

OpenStudy (raden):

let's check together :) see this identity : sin^2 (x) + cos^2 (x) = 1 cos^2 (x) = 1 - sin^2 (x) cos(x) = sqrt (1 - sin^2 (x)) now for evaluate cos(sin^-1(2x)), it can be cos(sin^-1(2x)) = sqrt (1 - sin^2 (sin^-1(2x)) ) = sqrt (1 - (2x)^2 ) = sqrt(1 - 4x^2)

OpenStudy (anonymous):

(4x)sqrt(1-4x^2)

OpenStudy (raden):

yes, totally is right

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (raden):

you're welcome ;)